OpenStudy (anonymous):

can someone verify the following: ((sin2x)/cot(x))=1-cot(2x)

OpenStudy (aum):

\(\sin^2(x)\) or \(\sin(2x)\) ?

OpenStudy (anonymous):

sin(2x)

OpenStudy (aum):

The right hand side is cot(2x) ?

OpenStudy (anonymous):

\[\frac{ 1 }{ cotx } = tanx\]

OpenStudy (anonymous):

the right side is 1-cot(2x)

OpenStudy (anonymous):

\[\sin(2x)\tan(x)=1-\cot(2x) \implies -1+\cot(2x)+\sin(2x)\tan(x)=0\]

OpenStudy (anonymous):

Well you can sub y = tan(x/2) but I have to go right now, c ya and good luck, if you still need help later I'll get back to you then.

OpenStudy (aum):

You can try substituting the following: sin(2x) = 2sin(x)cos(x) cot(x) = cos(x) / sin(x) cot(2x) = 1 / tan(2x) where tan(2x) = 2tan(x) / (1 - tan^2(x)); tan(x) = sin(x)/cos(x).

OpenStudy (anonymous):